# GCP: Guarded Collaborative Perception with Spatial-Temporal Aware Malicious Agent Detection

### v1

@@ Abstract
Collaborative perception significantly enhances autonomous driving safety by extending each vehicle’s perception range through message sharing among connected and autonomous vehicles. Unfortunately, it is also vulnerable to adversarial message attacks from malicious agents, resulting in severe performance degradation. While existing defenses employ hypothesis-and-verification frameworks to detect malicious agents based on single-shot outliers, they overlook temporal message correlations, which can be circumvented by subtle yet harmful perturbations in model input and output spaces. This paper reveals a novel blind area confusion (BAC) attack that compromises existing single-shot outlier-based detection methods. As a countermeasure, we propose GCP, a Guarded Collaborative Perception framework based on spatial-temporal aware malicious agent detection, which maintains single-shot spatial consistency through a confidence-scaled spatial concordance loss, while simultaneously examining temporal anomalies by reconstructing historical bird’s eye view motion flows in low-confidence regions. We also employ a joint spatial-temporal Benjamini-Hochberg test to synthesize dual-domain anomaly results for reliable malicious agent detection. Extensive experiments demonstrate GCP’s superior performance under diverse attack scenarios, achieving up to 34.69% improvements in AP@0.5 compared to the state-of-the-art CP defense strategies under BAC attacks, while maintaining consistent 5-8% improvements under other typical attacks. Code will be released at https://github.com/CP-Security/GCP.git.
@@ I Introduction
Collaborative perception (CP) supersedes single-agent perception by enabling information-sharing among multiple connected and autonomous vehicles (CAVs), substantially enlarging a vehicle’s perception scope and accuracy [1, 2, 3, 4, 5, 6, 7]. The extended perception helps an ego vehicle detect occluded objects that were originally difficult to recognize with single-vehicle perception due to physical occlusions, thereby boosting the safety of autonomous driving [8, 9, 10, 11, 12]. To collaborate, one simple way for an ego CAV to receive helps is to directly request early-stage raw data or late-stage detection results from the neighboring CAVs, and then combine this information with its own data to get the CP results. However, these methods are either bandwidth-consuming or vulnerable to perception noise or malicious message attacks. The recent development of deep learning has facilitated feature-level fusion, where the collaborative CAVs send an intermediate representation of deep neural network models to an ego CAV for aggregation, enhancing the performance-bandwidth trade-off in multi-agent perception.
@@
To guard CP systems, most existing works adopt outlier-based defense against malicious agents following a multi-round hypothesize-and-verify paradigm. In each round, ego agents first generate multiple hypothetical CP outcomes by assuming portion of the collaborators are benign based on prior knowledge and then verify the consistency between the ego agent’s single perception outcome and CP results. The iteration repeats until all malicious agents are identified. Following this idea, Li et al. [15] developed ROBOSAC, a random sample consensus-based method for detecting malicious agents. Besides, Zhao et al. [16] proposed MADE, a multi-test detection framework utilizing match and reconstruction losses to gauge consensus between the ego CAV and collaborative results. Despite its efficacy, the existing hypothesize-and-verify-based CP defense method merely compares the spatial outlier consistency in a single frame without considering the temporal contexts between consecutive frames. Meanwhile, the static outlier-based detection can be easily bypassed when the attacks are subtle yet dangerous in both model input and output. CP-Guard+ [17] proposed a feature-level defense by directly comparing the divergence of the spatial bird eye’s view (BEV) feature map of the ego agent and collaborators, effectively increasing the system’s scalability. However, all the above works only leverage independent spatial-domain information in a single time slot for malicious agent detection, without considering the temporal correlation of collaborators’ messages in different time slots, resulting in suboptimal defensive CP performance.
We strongly believe incorporating temporal knowledge is crucial for CP defense based on two key insights. First, adversarial attacks in real-world scenarios often exhibit distinct temporal patterns. When malicious agents inject perturbations intermittently to maintain stealth, these attacks manifest as anomalous variations in the temporal domain of CP results. For instance, both ROBOSAC [15] and MADE [16] observe that attackers typically alternate between sending malicious and benign messages across time slots to avoid detection. This temporal characteristic provides an additional verification dimension: beyond examining spatial consistency in the current frame, we can leverage historical clean messages from collaborators as reliable references to identify suspicious temporal deviations. Second, even when attackers continuously inject perturbations, their impact on CP results inevitably creates distinctive temporal patterns that differ from normal CP behaviors. These patterns manifest in various aspects, such as unnatural object motion trajectories, inconsistent detection confidence variations, or abrupt changes in spatial feature distributions across frames. Such temporal anomalies, while potentially subtle in individual frames, become more apparent when analyzed over extended long sequences.
@@
  • We conduct comprehensive experiments on diverse attack scenarios with V2X-Sim dataset. The results demonstrate that GCP achieves the state-of-the-art performance, with up to 34.69% improvements in AP@0.5 compared to the existing state-of-the-art defensive CP methods under intense BAC attacks, while maintaining 5-8% advantages under other typical adversarial attacks targeting at CP systems.
@@ II-A Collaborative Perception (CP)
To overcome the inherent limitations of single-agent perception systems, particularly their restricted field-of-view (FoV) and environmental occlusions, collaborative perception (CP) has emerged as a promising paradigm that leverages multi-agent data fusion to enhance perception comprehensiveness and accuracy [18, 19]. The evolution of CP systems has witnessed various fusion strategies, from early approaches like raw-data-level fusion [1] and output-level fusion [20], which faced significant challenges in communication overhead and perception quality due to their high bandwidth requirements and information loss, respectively, to more sophisticated intermediate-level feature fusion methods. Notable advances in intermediate fusion include DiscoNet [21], which employs a teacher-student framework to learn an optimized collaboration graph for efficient knowledge transfer, and V2VNet [22], which leverages graph neural networks for efficient neighboring vehicle information aggregation in dynamic traffic scenarios. Where2comm [4] further advances the field by introducing a confidence-aware attention mechanism that simultaneously optimizes communication efficiency and perception performance through selective information sharing. While these developments have improved performance-bandwidth trade-offs, they have exposed critical vulnerabilities in CP systems, particularly regarding robustness against adversarial attacks and system failures, which is the main focus of this paper.
@@ II-C Defensive CP
In response to emerging threats, the research community has developed various defensive strategies, primarily focusing on output-level malicious agent detection through hypothesis-and-verification frameworks, each offering unique approaches to system security. ROBOSAC [15] employs an iterative approach with Hungarian matching to achieve perception consensus among collaborators, systematically identifying and excluding potentially malicious agents from the collaboration process. MADE [16] enhances this concept by introducing a comprehensive multi-test framework that leverages both match loss and collaborative reconstruction loss to ensure robust consistency verification among perception results. Zhang et al. [13] further contribute to this effort by utilizing occupancy maps for discrepancy detection, providing an additional layer of security through spatial consistency checking. While these outlier-based detection methods show promise in identifying obvious attacks, they remain vulnerable to sophisticated attacks that produce subtle yet dangerous perturbations in CP outputs, particularly in scenarios involving coordinated malicious activities. Recent advances have begun exploring intermediate feature-level inconsistency detection to improve both efficiency and detection accuracy [17]. However, existing approaches generally overlook the temporal dimension of collaborative messages and assume simplified intrusion models, limiting their effectiveness against advanced persistent threats. To address these limitations, we propose a comprehensive defense framework that considers both spatial and temporal aspects of malicious agent detection, offering a more robust and complete solution to the security challenges in CP systems.
@@ III-B Adversarial Threat Model
The CP framework defined in Section III-A becomes vulnerable when malicious agents exist among the collaborative CAVs. At time $t_{i}$, we denote the set of malicious agents as $\mathcal{M}_{t_{i}}\subseteq\mathcal{N}$. We assume that malicious agents can request and transmit collaborative messages to benign agents before being identified, and their adversarial messages follow a specific time distribution $\mathcal{D}$ (detailed in Section V-A). Once the victim CAV recognizes the malicious agents at a certain frame, it will discard the received messages from those malicious agents and cut off the connections with them until the next frame starts. Meanwhile, the malicious agents are assumed to have white-box access [25] to the model parameters since all agents in the CP system must share the same model architecture to enable feature-level cooperation. The malicious agents attack the CP system by transmitting perturbed feature maps to the victim CAV, resulting in its perception degradation. A general attack procedure can be described as follows.
  1. 1.
    Observation Encoding: Each malicious agent encodes its raw observation at time $t_{j}$ into an initial feature map $\mathbf{F}^{t_{j}}_{j}=\Phi_{\mathtt{{enc}}}^{j}(\mathbf{O}_{j}^{t_{j}})$, preparing to combine with adversarial perturbations.
  2. 2.
    Perturbation Generation: In general cases, the malicious agents generate perturbations [26, 27] by multi-step iterations, aiming to maximize the distance between CP results and ground truth (GT) while ensuring the input perturbation amplitude is lower than a certain value (using local prediction if GT is not available), which is represented by:
    where $\delta$ is the perturbation, $\Delta_{i}$ is the maximum allowable input perturbation amplitude, $\mathcal{L}(\cdot)$ is the Loss function, $\mathbf{Y}_{\delta}^{t_{j}}$ and $\mathbf{Y}_{\mathtt{gt}}^{t_{j}}$ represent the perturbed CP results and the ground truth at time $t_{j}$, respectively. However, single-shot outlier-based detection can easily identify this general perturbation generation. Section III-C introduces a more elaborated attack targeted at CP systems.
  3. 3.
    Attack Triggering: After multi-step iterations, the malicious agents generate intricately optimized perturbation and send it to the victim agent. Without defense, the victim agent will incorporate this adversarial message into the fusion based on Eq. 1, yielding attacked CP results with low accuracy.
@@ III-C Blind Area Confusion Attack
where $Dist(s,e)$ computes the Euclidean distance between grid $s$ and victim grid $e$, $D_{norm}=\sqrt{H^{2}+W^{2}}$ is the normalization factor based on BEV detection map dimensions (height $H$ and width $W$), $K_{base}=6$ establishes a hexagonal-like growth pattern, and $\gamma_{d}=0.3$ controls the decay rate. This distance-adaptive design ensures denser expansion near the victim and sparser expansion in distant regions, naturally modeling spatial perception reliability [28]. The algorithm expands both regions simultaneously through $Queue.Pop()$ (which removes and returns the next grid point from the front of the queue) and $Queue.Append()$ (which adds a new grid point to the end of the queue) operations. The initial queues $\mathbf{Q}_{ca}$ and $\mathbf{Q}_{ba}$ are created using $Queue\{m\}$, which constructs priority queues containing all grid points $m$ that are labeled as confident area (1) or blind area (-1) respectively. These queue operations enable systematic breadth-first expansion of regions. As grid points are popped from either queue for processing, their unassigned neighbors receive the corresponding label (1 for CA, -1 for BA). They are appended to the appropriate queue for future expansion. The final binary mask $\mathbf{M}_{c}^{t_{k}}$ is created by converting BA labels (-1) to 0, yielding a binary representation where 1 indicates confident areas and 0 represents blind areas. This region-growing process naturally forms boundaries between different perception zones, effectively capturing the spatial characteristics of the victim’s perception capabilities.
@@
where $\mathcal{L}(\cdot)$ denotes the optimization loss function, $\mathbf{M}_{\delta}=\mathbf{1}-\mathbf{M}_{c}^{t_{k}}\{\mathbf{D}_{i}^{t_{k}}\{\mathbf{Y}_{\delta}^{t_{k}}\}\}$ is the inverted confidence mask, $\sigma(\cdot)$ is the sigmoid activation function, ||$\cdot$|| ⋅ | computes element-wise absolute values, $\Delta_{i}$ bounds the input perturbation magnitude, $\Delta_{o}$ bounds the output perturbation magnitude, and $w_{\delta}$ is a positive weighting parameter.
@@
Algorithm 1 Blind Region Segmentation (BRS)
Input:
  • $\mathbf{M}_{c}^{t_{k}}$, initial confidence mask (all-zero matrix).
  • $\mathbf{Y}_{vic}^{t_{k}}$, detected bounding boxes by victim vehicle.
  • $\mathbf{Y}_{nvic}^{t_{k}}$, undetected bounding boxes by victim vehicle.
  • $\mathbf{D}_{i}^{t_{k}}$, the BEV detection map of CAV $i$.
Output: $\mathbf{M}_{c}^{t_{k}}$, binary confidence mask.
1:procedure BRS($\mathbf{D}_{i}^{t_{k}}$)
2:     $e\leftarrow Cluster(\mathbf{Y}_{vic}^{t_{k}})$triangleright Victim Grid
3:     if $\mathbf{Y}_{nvic}^{t_{k}}=\emptyset$ then
5:     end if
6:     $\mathbf{M}_{c}^{t_{k}}\{\mathbf{D}_{i}^{t_{k}}\{\mathbf{Y}_{vic}^{t_{k}}\}\}\leftarrow 1$
7:     $\mathbf{M}_{c}^{t_{k}}\{\mathbf{D}_{i}^{t_{k}}\{\mathbf{Y}_{nvic}^{t_{k}}\}\}\leftarrow-1$
8:     $\mathbf{Q}_{ca},\mathbf{Q}_{ba}\leftarrow Queue\{m\},\forall\ \mathbf{M}_{c}^{t_{k}}\{m\}=1\ \text{or}\ -1$
9:     while $\mathbf{Q}_{ca}\neq\emptyset$ or $\mathbf{Q}_{ba}\neq\emptyset$ do $\triangleright$ Region Growth
10:         for each $\mathbf{Q}\in\{\mathbf{Q}_{ca},\mathbf{Q}_{ba}\}$ and $\mathbf{Q}\neq\emptyset$:
11:          $s\leftarrow\mathbf{Q}.Pop()$
12:          for $j\in GetNeighbors(s,e)$ and $\mathbf{M}_{c}^{t_{k}}\{j\}=0$:
13:           $\mathbf{M}_{c}^{t_{k}}\{j\}\leftarrow[\mathbf{Q}=\mathbf{Q}_{ca}]?1:-1$
14:           $\mathbf{Q}.Append(j)$
15:     end while
16:     $\mathbf{M}_{c}^{t_{k}}\{m\}\leftarrow 0,\forall\ \mathbf{M}_{c}^{t_{k}}\{m\}=-1$triangleright Binary Mask
17:     return $\mathbf{M}_{c}^{t_{k}}$
18:end procedure

### v2

@@ Abstract
Collaborative perception significantly enhances autonomous driving safety by extending each vehicle’s perception range through message sharing among connected and autonomous vehicles. Unfortunately, it is also vulnerable to adversarial message attacks from malicious agents, resulting in severe performance degradation. While existing defenses employ hypothesis-and-verification frameworks to detect malicious agents based on single-shot outliers, they overlook temporal message correlations, which can be circumvented by subtle yet harmful perturbations in model input and output spaces. This paper reveals a novel blind area confusion (BAC) attack that compromises existing single-shot outlier-based detection methods. As a countermeasure, we propose GCP, a Guarded Collaborative Perception framework based on spatial-temporal aware malicious agent detection, which maintains single-shot spatial consistency through a confidence-scaled spatial concordance loss, while simultaneously examining temporal anomalies by reconstructing historical bird’s eye view motion flows in low-confidence regions. We also employ a joint spatial-temporal Benjamini-Hochberg test to synthesize dual-domain anomaly results for reliable malicious agent detection. Extensive experiments demonstrate GCP’s superior performance under diverse attack scenarios, achieving up to 34.69% improvements in AP@0.5 compared to the state-of-the-art CP defense strategies under BAC attacks, while maintaining consistent 5-8% improvements under other typical attacks. Code will be released at https://github.com/yihangtao/GCP.git.
@@ I Introduction
Collaborative perception (CP) supersedes single-agent perception by enabling information-sharing among multiple connected and autonomous vehicles (CAVs), substantially enlarging a vehicle’s perception scope and accuracy [5, 18, 11, 15, 12, 7, 25, 31]. The extended perception helps an ego vehicle detect occluded objects that were originally difficult to recognize with single-vehicle perception due to physical occlusions, thereby boosting the safety of autonomous driving. To collaborate, one simple way for an ego CAV to receive helps is to directly request early-stage raw data or late-stage detection results from the neighboring CAVs, and then combine this information with its own data to get the CP results. However, these methods are either bandwidth-consuming or vulnerable to perception noise or malicious message attacks. The recent development of deep learning has facilitated feature-level fusion, where the collaborative CAVs send an intermediate representation of deep neural network models to an ego CAV for aggregation, enhancing the performance-bandwidth trade-off in multi-agent perception.
@@
Existing CP defense methods are primarily unsupervised outlier detection schemes based on the hypothesize-and-verify paradigm. For instance, ROBOSAC [17] and CP-Guard [14] employ sample consensus strategies to identify malicious agents, while MADE [35] utilizes multi-test consistency checks. Recently, CP-Guard+ [2] utilizes prior knowledge of attack patterns for training, which accelerates the detection speed yet imposes stronger assumption on the defender’s knowledge. Furthermore, all these methods rely heavily on spatial information within a single slot, ignoring temporal correlations across frames. Consequently, they remain vulnerable to dynamic attacks that are spatially subtle but temporally anomalous.
We strongly believe incorporating temporal knowledge is crucial for CP defense based on two key insights. ❶ First, adversarial attacks in real-world scenarios often exhibit distinct temporal patterns. When malicious agents inject perturbations intermittently to maintain stealth, these attacks manifest as anomalous variations in the temporal domain of CP results. For instance, both ROBOSAC [17] and MADE [35] observe that attackers typically alternate between sending malicious and benign messages across time slots to avoid detection. This temporal characteristic provides an additional verification dimension: beyond examining spatial consistency in the current frame, we can leverage historical clean messages from collaborators as reliable references to identify suspicious temporal deviations. ❷ Second, even when attackers continuously inject perturbations, their impact on CP results inevitably creates distinctive temporal patterns that differ from normal CP behaviors. These patterns manifest in various aspects, such as unnatural object motion trajectories, inconsistent detection confidence variations, or abrupt changes in spatial feature distributions across frames. Such temporal anomalies, while potentially subtle in individual frames, become more apparent when analyzed over extended long sequences.
@@
  • We conduct comprehensive experiments on diverse attack scenarios with V2X-Sim [18] dataset. The results demonstrate that GCP achieves the state-of-the-art performance, with up to 34.69% improvements in AP@0.5 compared to the existing state-of-the-art defensive CP methods under intense BAC attacks, while maintaining 5-8% advantages under other adversarial attacks targeting at CP systems.
@@ II-A Collaborative Perception (CP)
To overcome the inherent limitations of single-agent perception, particularly restricted field-of-view (FoV) and occlusions, collaborative perception (CP) has emerged as a promising paradigm leveraging multi-agent fusion to enhance perception accuracy [10, 13]. The evolution of CP systems has witnessed various strategies, from early raw-data [5] and output-level fusion [33], which faced challenges in communication overhead and information loss, to sophisticated intermediate-level feature fusion. Notable advances include DiscoNet [19], which employs a teacher-student framework to learn an optimized collaboration graph, V2VNet [27], which leverages graph neural networks for efficient aggregation, and When2com [21], which learns to construct communication groups to decide when to communicate. Where2comm [15] further advances the field by introducing a confidence-aware attention mechanism that simultaneously optimizes communication efficiency and perception performance. Similarly, R-ACP [8] proposes a robust task-oriented communication strategy that optimizes online self-calibration and efficient feature sharing by minimizing the Age of Perceived Targets (AoPT) to ensure timely and accurate perception. While improving performance-bandwidth trade-offs, these developments exposed critical vulnerabilities in CP systems regarding robustness against adversarial attacks, which is the main focus of this paper.
@@ II-C Defensive CP
In response to emerging threats, the research community has developed various defensive strategies, primarily focusing on output-level malicious agent detection through hypothesis-and-verification frameworks. ROBOSAC [17] and CP-Guard [14] employ sample consensus strategies to systematically identify and exclude potentially malicious agents. MADE [35] utilizes a multi-test framework leveraging both match loss and collaborative reconstruction loss to ensure robust consistency. Zhang et al. [34] further utilize occupancy maps for discrepancy detection. While these unsupervised outlier-based detection methods show promise, they remain vulnerable to sophisticated attacks with subtle perturbations. Recently, CP-Guard+ [2] trains a feature classifier by pre-simulating attack patterns (CP-GuardBench). However, this approach relies on the strong assumption that the defender possesses prior knowledge of the attacker’s perturbation mechanism to construct the training dataset. This contrasts with other methods that operate without such prior knowledge or access to simulated attack data. Furthermore, existing approaches generally overlook the temporal information, limiting their effectiveness against dynamic threats. To address these limitations, we propose a comprehensive defense framework that considers both spatial and temporal aspects, offering a more robust solution to CP security challenges.
@@ III-A Model of CP
However, this collaborative process is vulnerable when malicious agents $\mathcal{M}_{t_{i}}\subseteq\mathcal{N}$ exist. These agents can encode raw observations into initial feature maps and then generate adversarial perturbations $\delta$ by maximizing the distance between CP results and ground truth (GT):
where $\mathcal{L}$ is the perception loss function, $\mathbf{Y}_{\delta}^{t_{j}}$ and $\mathbf{Y}_{\mathtt{gt}}^{t_{j}}$ denote the perturbed CP output and the ground truth at time $t_{j}$, respectively, and $\Delta_{i}$ is the maximum allowable perturbation amplitude. Finally, they transmit these perturbed features to the victim to degrade perception accuracy.
@@ III-B Adversarial Threat Model
We adopt the same threat model assumptions as in previous works [17, 35, 14]. Specifically, we consider an internal attacker who has compromised at least one legitimate CAV and gained white-box access to the shared CP model architecture, parameters, and the feature maps transmitted among vehicles. This access is inherent to the compromised vehicle’s legitimate participation in the CP system. The attacker aims to mislead the victim’s perception by sending malicious feature maps while maintaining stealthiness. We also assume the attacker adheres to the system’s communication protocol regarding timing and does not manipulate artificial delays or jitter, focusing instead on manipulating the semantic content of the transmitted features.
@@
Algorithm 1 Blind Region Segmentation (BRS)
Input: $\mathbf{M}_{c}^{t_{k}}$ (initial confidence mask), $\mathbf{Y}_{vic}^{t_{k}}$ (victim detections), $\mathbf{Y}_{nvic}^{t_{k}}$ (unseen detections), $\mathbf{D}_{i}^{t_{k}}$ (BEV detection map)
Output: $\mathbf{M}_{c}^{t_{k}}$ (binary confidence mask)
1:procedure BRS($\mathbf{D}_{i}^{t_{k}}$)
2:  $e\leftarrow Cluster(\mathbf{Y}_{vic}^{t_{k}})$ $\triangleright$ Victim Grid
3:  if $\mathbf{Y}_{nvic}^{t_{k}}=\emptyset$ then
5:  end if
6:  $\mathbf{M}_{c}^{t_{k}}\{\mathbf{D}_{i}^{t_{k}}\{\mathbf{Y}_{vic}^{t_{k}}\}\}\leftarrow 1$; $\mathbf{M}_{c}^{t_{k}}\{\mathbf{D}_{i}^{t_{k}}\{\mathbf{Y}_{nvic}^{t_{k}}\}\}\leftarrow-1$
7:  $\mathbf{Q}_{ca},\mathbf{Q}_{ba}\leftarrow Queue\{m\},\forall\ \mathbf{M}_{c}^{t_{k}}\{m\}=1\ \text{or}\ -1$
8:  while $\mathbf{Q}_{ca}\neq\emptyset$ or $\mathbf{Q}_{ba}\neq\emptyset$ do $\triangleright$ Region Growing
9:    for $\mathbf{Q}\in\{\mathbf{Q}_{ca},\mathbf{Q}_{ba}\}$ and $\mathbf{Q}\neq\emptyset$ do
10:     $s\leftarrow\mathbf{Q}.Pop()$
11:     for $j\in GetNeighbors(s,e)$ and $\mathbf{M}_{c}^{t_{k}}\{j\}=0$ do
12:      $\mathbf{M}_{c}^{t_{k}}\{j\}\leftarrow[\mathbf{Q}=\mathbf{Q}_{ca}]?1:-1$
13:      $\mathbf{Q}.Append(j)$
14:     end for
15:    end for
16:  end while
17:  $\mathbf{M}_{c}^{t_{k}}\{m\}\leftarrow 0,\forall\ \mathbf{M}_{c}^{t_{k}}\{m\}=-1$ $\triangleright$ Binary Mask
18:  return $\mathbf{M}_{c}^{t_{k}}$
19:end procedure
@@ III-C Blind Area Confusion Attack
where $Dist(s,e)$ computes the Euclidean distance between grid $s$ and victim grid $e$, $D_{norm}=\sqrt{H^{2}+W^{2}}$ is the normalization factor based on BEV detection map dimensions (height $H$ and width $W$), $K_{base}=6$ establishes a hexagonal-like growth pattern, and $\gamma_{d}=0.3$ controls the decay rate. This distance-adaptive design ensures denser expansion near the victim and sparser expansion in distant regions, naturally modeling spatial perception reliability [32]. The expansion proceeds simultaneously via queue operations. Initialized by $Queue\{m\}$, priority queues $\mathbf{Q}_{ca}$ and $\mathbf{Q}_{ba}$ drive breadth-first growth. Grids are iteratively processed using $Queue.Pop()$, where unassigned neighbors inherit the label (1 for CA, -1 for BA) and are added via $Queue.Append()$. Finally, converting BA labels to 0 generates the binary mask $\mathbf{M}_{c}^{t_{k}}$, which effectively delineates the victim’s perception boundaries.
@@
where $\mathcal{L}(\cdot)$ denotes the optimization loss function, $\mathbf{M}_{\delta}=\mathbf{1}-\mathbf{M}_{c}^{t_{k}}\{\mathbf{D}_{i}^{t_{k}}\{\mathbf{Y}_{\delta}^{t_{k}}\}\}$ is the inverted confidence mask, $\sigma(\cdot)$ is the sigmoid activation function, $|\cdot|$ computes element-wise absolute values, $\Delta_{i}$ bounds the input perturbation magnitude, $\Delta_{o}$ bounds the output perturbation magnitude, and $w_{\delta}$ is a positive weighting parameter.